OpenStudy (anonymous):

What property is shown by the equation?–10 + (–15) = –15 + (–10)

OpenStudy (anonymous):

@shinalcantara

OpenStudy (anonymous):

\[-10+(-15)=-15+(-10)\]

OpenStudy (shinalcantara):

Commutative property of addition

OpenStudy (shinalcantara):

that property simply implies that a + b = b + a